A liquid -in - glass thermometer has a bulb of 4 cm3 and the bore of the tube is 0.2 cm in diameter. The interval between the fixed points is 10 cm. If the thermometer is heated from 0oC to 1000C, calculate the apparent expansivity of the liquid
1
Expert's answer
2020-06-12T11:11:58-0400

We know that between 0 and 10 cm the temperature range is 0 to 100°C. Therefore, the expansivity is


"\gamma=\frac{\Delta H/H}{\Delta t}=\frac{\Delta H/[V/(\pi r^2)]}{\Delta t}=\frac{\pi d^2\Delta H}{4V\Delta t},\\\space\\
\gamma=\frac{3.14\cdot0.2^2\cdot0.1}{4\cdot4\cdot100}=7.9\cdot10^{-6}\text{/°C}."